# ToolRM: Outcome Reward Models for Tool-Calling Large Language Models

### v1

@@ Abstract
As large language models (LLMs) increasingly interact with external tools, reward modeling for tool use has become a critical yet underexplored area. Existing reward models, trained primarily on natural language outputs, struggle to evaluate tool-based reasoning and execution. To quantify this gap, we introduce FC-RewardBench, the first benchmark designed to systematically assess reward models’ performance in tool-calling scenarios. Our analysis shows that current reward models often miss key signals of effective tool use, highlighting the need for domain-specific modeling. To address this, we propose a training framework for outcome-based reward models using data synthesized from permissively licensed, open-weight LLMs. We train models ranging from 1.7B to 14B parameters and evaluate them across seven out-of-domain benchmarks. These models consistently outperform general-purpose baselines, achieving up to 25% average improvement in downstream task performance and enabling data-efficient fine-tuning through reward-guided filtering.
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 Achiam et al. (2023), Claude, and Gemini Reid et al. (2024) have rapidly advanced the field of artificial intelligence (AI), demonstrating impressive performance across a wide range of natural language processing tasks. These models are capable of generating coherent text, answering complex questions, writing code, and even reasoning over multi-step problems. As LLMs become increasingly central to AI systems deployed in real-world settings, the demand for them to go beyond language generation and interact with external systems and tools has grown substantially. To meet this demand, tool calling has emerged as a critical capability that allows LLMs to invoke external functions — such as APIs, databases, calculators, and search engines Prabhakar et al. (2025); Zhang et al. (2024); Abdelaziz et al. (2024); Liu et al. (2024b); Lin et al. (2024). This paradigm shifts the role of LLMs from standalone generators of text to orchestrators of complex workflows, making them more reliable and useful in practical applications such as autonomous agents, virtual assistants, and multi-modal systems.
@@
To effectively train and fine-tune LLMs, reward models are a core component of reinforcement learning (RL) and preference optimization pipelines Wang et al. (2023). They provide a learned signal that estimates output quality, enabling scalable evaluation without requiring human judgment on every example. Broadly, reward models fall into two categories: process reward models (PRMs) Lightman et al. (2023), which score intermediate reasoning steps, and outcome reward models (ORMs) Cobbe et al. (2021), which only evaluate the final answer. Process models offer finer control over reasoning but demand costly, fine-grained annotations. In contrast, outcome models are easier to train and can achieve comparable performance gains Uesato et al. (2022). Because outcome feedback is naturally aligned with how users assess responses (e.g., “this answer is helpful”), it is cheaper to collect and more scalable. While outcome models do not guarantee step-by-step reliability, they are often sufficient to guide LLMs toward high-quality outputs, particularly in tasks where intermediate reasoning is difficult to evaluate.
@@
Figure 1: Performance of ToolRM, top reward models from RewardBench, and leading LLMs-as-judges on FC-RewardBench. Note: We abbreviate the model names for conciseness, for example, L3.1-$x$ corresponds to Llama-3.1-$x$; SW-Rew-$x$ corresponds to SkyWorks-Reward-$x$. Full model names can be found in the Appendix.
Reward modeling for tool calling remains an underexplored area. Existing reward models are typically trained on natural language outputs and are not well-suited to capture the nuances of tool-based reasoning and execution. Currently, there is no dedicated benchmark to systematically evaluate the effectiveness of reward models in tool-calling scenarios. To address this gap, first, we introduce FC-RewardBench [^1] https://huggingface.co/datasets/ibm-research/fc-reward-bench – a comprehensive benchmark specifically designed to evaluate reward models on tool-calling tasks. Derived from the Berkeley Function Calling Leaderboard (BFCL) Version 3 Patil et al. (2025), the dataset consists of 1500 unique user inputs along with both correct and incorrect function calls. Second, we apply several state-of-the-art general-purpose reward models trained on human preferences for natural language generation to our benchmark (Figure 1). Our analysis reveals that these models often fail to capture key aspects of successful tool use, highlighting the need for domain-specific reward modeling in this area.
To address this issue, we introduce ToolRM - a collection of specialized ORMs for tool calling. Trained on preference data synthesized from a diverse set of open-source function-calling models, ToolRM outperforms much larger reward models and LLMs-as-Judges on FC-RewardBench. In downstream applications, ToolRM demonstrates up to 25% average improvement across multiple benchmarks in a Best-of-$n$ setting. These models also enable efficient data filtering, yielding better fine-tuned models with less data.
@@
  • We introduce FC-RewardBench; the first benchmark to evaluate reward models on a tool calling setting with strong correlation to downstream task performance.
  • We propose a framework for training a tool calling ORM using data generated from permissively-licensed, medium-sized, open-weight LLMs.
  • We train multiple reward models (ToolRM) varying in size from 1.7B to 14B in parameters, and extensively evaluate our proposed models on seven out-of-domain benchmarks.
  • We show that our reward models lead to up to 25% improvement on average across all evaluated benchmarks in a Best-of-$n$ setting. We also show that our reward models can be utilized for data filtering and lead to finetuned models that are better in performance with significantly less data.
@@ 2.2 RL for Tool-Use Alignment
Reinforcement Learning (RL) has recently emerged as a powerful approach for aligning LLMs with effective tool use. Search-R1 Jin et al. (2025) trains LLMs to iteratively generate and refine search queries during multi-step reasoning, showing that RL feedback helps balance exploration and precision in retrieval. ToRL Li et al. (2025) enables base models to autonomously discover tool-use strategies, with reward signals driving emergent behaviors such as strategic invocation and adaptive switching between reasoning modes. ReTool Feng et al. (2025) further interleaves code execution with natural language reasoning, using outcome feedback to guide when and how tools are invoked, yielding strong gains in mathematical problem solving.
Several works explore reward design for tool use. ToolRL Qian et al. (2025a) provides a systematic study, highlighting how reward type, granularity, and temporal dynamics impact alignment. StepTool Yu et al. (2024) applies step-level reward shaping and policy-gradient optimization, improving performance on complex multi-step tasks. CodeTool Lu et al. (2025) combines RL with step-level supervision, encouraging models to reason about intermediate states. SWE-RL Wei et al. (2025) leverages large-scale software evolution data to optimize reasoning over evolving sequences of actions, demonstrating how RL can capture temporal dependencies in tool use. iTool Zeng et al. (2025) addresses performance decay from synthetic data through iterative reinforced fine-tuning with Monte Carlo Tree Search, enhancing robustness in complex scenarios.
Collectively, these works demonstrate the effectiveness of RL in aligning LLMs toward general-purpose tool use. However, to the best of our knowledge, none of these methods explicitly employs an ORM that directly evaluates or optimizes the overall quality of an entire sequence of tool interactions.
@@ 2.3 Reward Modeling
Reward models (RMs) provide scalar preference signals that guide LLMs through preference optimization or RL, and can be broadly divided into ORMs (outcome reward models), which only evaluate the final output, and PRMs (process reward models), which score intermediate reasoning steps. Early verifier-based approaches in the math domain Cobbe et al. (2021) laid the foundation for ORMs, while later work explicitly contrasted outcome- and process-based supervision for math word problems Uesato et al. (2022), and developed PRMs that reward coherent stepwise reasoning Lightman et al. (2023). Despite their promise, PRMs often face robustness and supervision challenges, as highlighted by failed attempts reported in the DeepSeek-R1 work Guo et al. (2025). In contrast, ORMs have proven more scalable, focusing on final correctness and generalization, with recent advances such as Skywork-Reward Liu et al. (2024a) demonstrating effective recipes for outcome-based training that achieved state-of-the-art performance on RewardBench Lambert et al. (2024). While prior work has studied RMs primarily in free-text reasoning and math/code domains, to the best of our knowledge, this is the first work to introduce ORMs for tool calling, where outcomes are defined over tool call sequences.
@@ 3.1 FC-RewardBench Evaluation Dataset
Figure 2: Representative example from the FC-RewardBench dataset, where the parameter player_count is set to an incorrect value. The tool catalog is hidden for brevity.
@@
While several benchmarks have been proposed to evaluate RMs on tasks involving chat, reasoning, safety (Lambert et al., 2024); factuality, instruction following, math (Malik et al., 2025); sensitivity to style and subtle content differences (Liu et al., ); and multi-modal judges for image-generation models (Chen et al., ), there remains a notable gap in the evaluation of RMs for function-calling tasks. To bridge this gap, we propose FC-RewardBench, a benchmark dataset specifically designed to evaluate RMs on function-calling tasks. This dataset comprises 1500 unique data points, each containing a tool catalog, a user query, and the associated correct and incorrect tool calls for a given user query.
To construct FC-RewardBench, we utilize the single-turn splits of the BFCL-v3 dataset (Yan et al., 2024). The tool catalog, user query, and the correct tool calls in the dataset are directly sourced from BFCL-v3. Incorrect tool calls are generated using a pool of 25 permissively licensed models, spanning sizes from 0.5B to 685B parameters. Each model is prompted to generate a tool call in response to the user query. The outputs are compared against the ground-truth tool calls, and only the incorrect generations are retained. From this pool, we randomly sample one incorrect call per instance to prevent over-representation from any single user query. Finally, 1,500 such examples are randomly selected to form the final dataset.
Figure 2 shows a representative example from the dataset, where the incorrect tool call sets the player_count parameter to an incorrect value of 5 instead of the correct value 6. Table 1 presents a breakdown of error types observed in the dataset. Notably, a majority of the incorrect calls involve subtle errors such as incorrect parameter values, missing optional parameters, or an incorrect number of functions, which are non-trivial to detect. These characteristics require the RM to demonstrate a deeper understanding of the function-calling task, making FC-RewardBench a challenging and discriminative benchmark. Additional details about the benchmark are provided in Appendix A.1.
3.2 ORM Training Data Generation
To train RMs for function-calling tasks, we need data that contains the following fields: (a) a user query, (b) a tool catalog – a list of tools available to the model to answer the user query, (c) the corresponding correct tool call, and (d) one or more incorrect tool calls. To generate this training data, we leverage a diverse pool of open-source, permissively licensed, medium-sized models with function-calling capabilities.
Using publicly available function-calling datasets, which include user queries, tool catalogs, and corresponding ground-truth tool call sequences, we prompt each model to complete the task. Specifically, each model is instructed to generate a tool call in response to a user query using the tools provided in the dataset. This process simulates a real-world inference scenario where the model must decide which tool to invoke and how to structure the call.
Once the model-generated tool call is produced, we compare it against the ground-truth tool call specified in the dataset. If the model output deviates from the ground truth, it is retained as an example of an incorrect tool call. Conversely, outputs that exactly match the ground truth are discarded, as they do not contribute to the error-focused training objective.
This data generation approach captures the nuance and variability of real-world model behavior, reflecting the kinds of errors that actual models tend to make in practice, including subtle and complex failure modes that may be difficult to anticipate or enumerate manually.
@@ 3.3 Reward Modeling
A RM aims to capture human preferences between two model outputs. This is typically modeled by the Bradley-Terry model Bradley and Terry (1952), which defines the probability that output $y_{+}$ is preferred over $y_{-}$, given an input $x$ as:
$p(y_{+}\succ y_{-}\mid x)=\\ \frac{\exp(r(x,y_{+}))}{\exp(r(x,y_{+}))+\exp(r(x,y_{-}))}$ (1)
where $r(x,y)$ is a scalar reward function assigning a relative score to output $y$ given input $x$.
Training requires curating a dataset of pairwise preferences $D=\{(x,y_{+},y_{-}):y_{+}\succ y_{-}\}$, with preferences being obtained through either human annotations Stiennon et al. (2020); Ouyang et al. (2022) or synthetic generation methods Pace et al. (2024); Hosseini et al. (2024). The reward function $r$ is parameterized by a neural network $r_{\theta}$, typically initialized from a supervised fine-tuned model with the final layer replaced by a linear head.
@@
where $\sigma$ denotes the sigmoid function.
In this work, we use reward centering Eisenstein et al. (2023) to ensure that rewards are zero-centered. This is achieved by adding the following regularization term to the optimization objective:
@@ 4 Experimental Setup
To create training data for the RM, we select open-source datasets that cover various aspects of function-calling, such as the API-Gen dataset Liu et al. (2024c) for single-turn interactions, the Schema-Guided Dialogue (SGD) dataset Rastogi et al. (2020) for multi-turn interactions with tool invocations and responses, and the xlam-irrelevance [^2] https://huggingface.co/datasets/MadeAgents/xlam-irrelevance-7.5k dataset for cases where the system lacks sufficient information to respond to a user query.
Since these datasets are common training datasets and our primary focus is to elicit representative incorrect behavior from the model, we follow Lin et al. (2024) and obfuscate the data samples to avoid the model regurgitating its training data. We obfuscate the samples by replacing function and parameter names with randomly generated strings and reordering the keys in the function schema. Note that this transformation preserves the semantics of the input–output pairs.
We then use a collection of 11 permissively-licensed, medium-sized, open-weight models to generate the training data. The pool includes both general-purpose instruction-tuned models with function-calling capabilities and function-calling specific models, with parameter counts ranging from 0.5B to 32B. Specifically, we use the IBM Granite 3.3-Instruct Granite Team and Qwen2.5-Instruct series Yang et al. (2025) models along with Granite-20b-function-calling Abdelaziz et al. (2024), SmolLM2 Allal et al. (2025), Mistral-7b-Instruct-v0.3333https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 and Mistral-Nemo-Instruct-2407444https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407.
After generating outputs from the model pool and keeping only the incorrect ones, we subsample one incorrect output per input user query to prevent over-representation from a user query in the training data. Overall, this results in 180,000 training data samples divided into 85,000 single and multi-turn data each, and 10,000 irrelevance data.
The full list of models used to generate the training data, along with a few training data samples, is provided in Appendix A.2.
@@
We use the Qwen-2.5-Instruct models Team (2024); Yang et al. (2024) as the base architecture for our RMs. Specifically, we select the 1.5B, 7B, and 14B parameter variants, as they are Apache-2.0 licensed and offer a practical balance between size and performance. We initialize the RMs with the instruction-tuned model weights and replace the final language modeling head with a linear layer that maps the hidden representation to a scalar reward value.
The RMs accept the specifications of available functions, conversation history, and the generated tool call as input and produce a scalar reward as output (refer to Appendix A.3 for prompt template). We train all RMs for 1 epoch with a learning rate set to 1e-6 and a cosine learning rate schedule with warmup set to 3% of total steps. We also use reward centering Eisenstein et al. (2023) with a coefficient set to 0.01.
@@
In addition to evaluating models’ performance on FC-RewardBench, we also evaluate on a variety of function calling benchmarks:
  • Berkeley Function-Calling Leaderboard (BFCL) Patil et al. (2025) is a comprehensive function calling leaderboard that includes a dataset of over 4,400 instances. The leaderboard evaluates model performance across single and multi-turn settings and include tasks that vary in difficulty and programming languages.
  • API-Bank (Li et al., 2023) has 314 tool-use dialogues with 753 API calls to assess LLMs’ capabilities in planning, retrieving, and calling APIs.
  • ToolAlpaca (Tang et al., 2023) test set has 271 tool-use instances spanning 50 distinct categories. Similar to Nexusflow.ai (2023), we used the simulated part of ToolAlpaca which has a total of 100 test examples.
  • NexusRaven API Evaluation [^5] https://huggingface.co/datasets/Nexusflow/NexusRaven_API_evaluation is a diverse benchmark for function calling with 318 test examples across 65 APIs, built by mining function definitions, docstrings, and execution contexts from open-source code, and
  • Sealtools Wu et al. (2024) is a large-scale self-instructed tool-learning dataset generated by ChatGPT, featuring diverse API-like tools with challenging nested, cross-field, and multi-tool calls.
@@
To evaluate performance on FC-RewardBench, we select eight RMs from RewardBench, spanning sizes from 3B to 70B parameters. We chose these models because they score highly on RewardBench and support tools in their chat template, which helps mitigate performance degradation due to prompt variability. In addition to these specialized RMs, we include six LLMs acting as judges, spanning sizes from 70B to 685B parameters.
To evaluate tool calling performance, we selected a wide range of models from BFCL V3 leaderboard [^6] https://gorilla.cs.berkeley.edu/leaderboard.html focusing on 1) the best performing models like xLAM-2777https://huggingface.co/collections/Salesforce/xlam-2-67ef5be12949d8dcdae354c4 (Llama-xLAM-2-70b-fc-r, xLAM-2-32b-fc-r, and Llama-xLAM-2-8b-fc-r), and Qwen3-32B [^8] https://huggingface.co/Qwen/Qwen3-32B and 2) small-sized models that perform well on the leaderboard to measure the extent of which our RM, ToolRM, can help, e.g. xLAM-2-1b-fc-r, xLAM-2-3b-fc-r, Qwen3-0.6B, Qwen3-1.7B, and Qwen3-4B.

### v2

@@ Abstract
As large language models (LLMs) increasingly interact with external tools, reward modeling for tool use has emerged as a critical yet underexplored area of research. Existing reward models, trained primarily on natural language outputs, struggle to evaluate tool-based reasoning and execution. To quantify this gap, we introduce FC-RewardBench, the first benchmark to systematically evaluate reward models in tool-calling scenarios. Our analysis shows that current reward models frequently miss key signals of effective tool use, highlighting the need for domain-specific modeling. We address this by proposing a training framework for outcome reward models using data synthesized from permissively licensed, open-weight LLMs. We introduce ToolRM – a suite of reward models for tool-use ranging from 1.7B to 14B parameters. Across diverse settings, these models consistently outperform general-purpose baselines. Notably, they achieve up to a 25% improvement with Best-of-$n$ sampling, while also improving robustness to input noise, enabling effective data filtering, and supporting RL-training of policy models.
@@ 1 Introduction
Large language models (LLMs) have rapidly advanced the field of artificial intelligence (AI), achieving strong performance across a wide range of tasks, including complex question answering, code generation, and multi-step reasoning (Li et al., 2025b). As these models are increasingly deployed in real-world systems, the need for them to interact with external tools has become critical. Tool calling enables LLMs to invoke external functions such as APIs, databases, calculators, and search engines (Prabhakar et al., 2025b; Zhang et al., 2024; Abdelaziz et al., 2024; Liu et al., 2024b; Lin et al., 2024), shifting their role from standalone text generators to orchestrators of complex workflows. This capability underpins their application in autonomous agents, virtual assistants, and multimodal systems.
@@
Training these LLMs effectively requires reward models, which are integrated into the learning pipeline through reinforcement learning (RL), preference optimization (Wang et al., 2023), and rejection sampling fine-tuning (Touvron et al., 2023; Team, 2024). Reward models provide learned signals that estimate output quality, enabling scalable evaluation without requiring human judgment on every example. Broadly, they fall into two categories: process reward models (PRMs) (Lightman et al., 2023), which score intermediate reasoning steps, and outcome reward models (ORMs) (Cobbe et al., 2021), which evaluate only the final answer.
@@
Figure 1: Performance of ToolRM, top reward models from RewardBench, Tool-augmented RM (Themis), and leading LLMs-as-judges on FC-RewardBench. Note: Model names are abbreviated for conciseness (e.g., L3.1-$x$, SR-$x$, and SC-$x$ correspond to Llama-3.1-$x$, SkyWorks-Reward-$x$, and SkyWorks-Critics-$x$, respectively). Full model names are provided in Appendix A.4.
Despite their successes, current reward models are designed primarily for natural language outputs (Zhong et al., 2025). Reward modeling for tool calling remains an underexplored area, with two notable gaps: (a) no dedicated benchmark exists for evaluating reward models in the function-calling domain [^1] Tool-use, tool-calling, and function-calling are used interchangeably throughout the paper, and (b) existing reward models fail to capture the nuances of tool-based reasoning and execution. In order to address these gaps, we first introduce FC-RewardBench – a comprehensive benchmark specifically designed to evaluate reward models on tool-calling tasks. Derived from the Berkeley Function Calling Leaderboard (BFCL) Version 3 (Patil et al., 2025), the dataset contains 1500 user inputs paired with correct and incorrect function calls. We benchmark several state-of-the-art general-purpose reward models on FC-RewardBench, and our analysis (Figure 1) shows that these models often fail to capture key aspects of successful tool use, hence failing to capture the nuances of tool-based reasoning and execution. To this end, next, we introduce ToolRM, a collection of specialized ORMs for tool calling. Trained on preference data synthesized from a diverse set of open-source function-calling models, ToolRM outperforms much larger reward models and LLMs-as-Judges on FC-RewardBench. In downstream applications, ToolRM achieves up to 25% average improvement across multiple benchmarks in a Best-of-$n$ setting, enables effective data filtering that produces stronger fine-tuned models with significantly less training data, and enable RL-training of policy models without requiring ground-truth labels for reward computation.
@@
  • We introduce FC-RewardBench ($\S$3.1), the first benchmark for evaluating reward models for tool-calling task, and demonstrate that existing RMs struggle in this domain ($\S$5.1).
  • We propose a framework for training outcome reward models for tool-calling tasks using data generated from permissively licensed, open-weight LLMs, and train a suite of reward models (ToolRM-1.7B, 7B, 14B) ($\S$4).
  • We show that ToolRM deliver substantial practical gains:  (1) up to 25% improvement when used for Best-of-$n$ sampling ($\S$5.2);  (2) superior fine-tuned models using only 50% of the data after ToolRM-based filtering ($\S$5.3.1); and  (3) RL-trained policies using ToolRM rewards match the performance of models trained with ground-truth–dependent rewards ($\S$5.3.2).
@@ 2.2 RL for Tool-Use Alignment
Reinforcement Learning has become a powerful approach for aligning LLMs with effective tool use. Search-R1 (Jin et al., 2025) trains LLMs to iteratively refine search queries, showing RL feedback balances exploration and retrieval precision. ToRL (Li et al., 2025a) enables autonomous discovery of tool-use strategies, with rewards driving emergent behaviors like strategic invocation and adaptive reasoning. ReTool (Feng et al., 2025) interleaves code execution with natural language reasoning, using outcome feedback to guide tool invocation, improving mathematical problem solving. Several works focus on reward design: ToolRL (Qian et al., 2025a) studies how reward type, granularity, and temporal dynamics affect alignment; StepTool (Yu et al., 2024) uses step-level reward shaping and policy-gradient optimization for multi-step tasks; CodeTool (Lu et al., 2025) combines RL with step-level supervision to encourage intermediate reasoning; SWE-RL (Wei et al., 2025) leverages software evolution data to optimize reasoning over action sequences, capturing temporal dependencies; and iTool (Zeng et al., 2025) mitigates performance decay from synthetic data via iterative reinforced fine-tuning with Monte Carlo Tree Search. Together, these works show RL’s effectiveness in aligning LLMs for general-purpose tool use, though none explicitly employ an ORM that evaluates or optimizes entire sequences of tool interactions.
@@ 2.3 Reward Modeling
Reward models (RMs) provide scalar preference signals that guide LLMs during preference optimization or RL (Wang et al., 2024). RMs can be broadly classified as Outcome Reward Models (ORMs), which score only the final output, and Process Reward Models (PRMs), which evaluate intermediate reasoning steps (Zhong et al., 2025). Early verifier-based methods in math (Cobbe et al., 2021) established ORMs, while later work contrasted outcome- vs. process-based supervision (Uesato et al., 2022) and developed PRMs that reward coherent stepwise reasoning (Lightman et al., 2023). However, PRMs face robustness and supervision challenges (Zhang et al., 2025), as evidenced by failed attempts reported by Guo et al. (2025). ORMs, in contrast, have scaled more reliably (Lin et al., 2025), with advances like Skywork-Reward (Liu et al., 2024a) achieving state-of-the-art results on RewardBench (Lambert et al., 2024). More recently, tool-augmented reward models (Li et al., 2024) allow RMs to use external tools for more accurate preference scoring. While prior work has focused on free-text, math, and code domains, to our knowledge this is the first to introduce ORMs for tool calling, where outcomes are defined by sequences of tool calls.
@@ 3.1 FC-RewardBench Evaluation Dataset
While several benchmarks evaluate RMs on tasks involving chat, reasoning, safety (Lambert et al., 2024); factuality, instruction following, and math (Malik et al., 2025), there remains a notable gap in the evaluation of RMs for function-calling tasks. To bridge this gap, we propose FC-RewardBench, a benchmark specifically designed to evaluate RMs on function-calling tasks. This dataset comprises 1500 unique data points, each containing a user query, a tool catalog (tools available to the model to answer the user query), and the associated correct and incorrect tool calls for a given user query.
To construct FC-RewardBench, we utilize the single-turn splits of the BFCL-v3 dataset (Patil et al., 2025). The tool catalog, user query, and the correct tool calls in the dataset are directly sourced from BFCL-v3. Incorrect tool calls are generated using a pool of 25 language models, spanning sizes from 0.5B to 685B parameters. Each model is prompted to generate a tool call in response to the user query. The outputs are compared against the ground-truth, and only the incorrect generations are retained. From this pool, we randomly sample one incorrect call per instance to prevent over-representation from any single user query. Finally, 1,500 such examples are randomly selected to form the final dataset.
@@
Table 1 presents a breakdown of error types observed in the dataset. Notably, a majority of the incorrect calls involve subtle errors such as incorrect parameter values, missing optional parameters, or an incorrect number of functions, which are non-trivial to detect. These characteristics require the RM to demonstrate a deeper understanding of the function-calling task, making FC-RewardBench a challenging and discriminative benchmark. Figure 3 shows a representative example from the dataset, and additional details about the benchmark are provided in Appendix A.1.
@@ 3.2 Reward Modeling
For pairwise preference modeling, RMs are commonly formulated using the Bradley–Terry model (Bradley and Terry, 1952), which defines the probability that output $y_{+}$ is preferred over $y_{-}$ given an input $x$ as:
$p(y_{+}\succ y_{-}\mid x)=\frac{e^{(r(x,y_{+}))}}{e^{(r(x,y_{+}))}+e^{(r(x,y_{-}))}}\\ =\sigma(r(x,y_{+})-r(x,y_{-}))$ (1)
where $r(x,y)$ is a scalar reward function, and $\sigma$ is the sigmoid function.
Training requires curating a dataset of pairwise preferences $D=\{(x,y_{+},y_{-}):y_{+}\succ y_{-}\}$, with preferences obtained through either human annotations (Stiennon et al., 2020; Ouyang et al., 2022) or synthetic generation methods (Pace et al., 2024; Hosseini et al., 2024). The reward function $r$ is parameterized by a neural network $r_{\theta}$, typically initialized from a supervised fine-tuned model with the final layer replaced by a linear head.
@@
In this work, we use reward centering (Eisenstein et al., 2023) to ensure that rewards are zero-centered. This is achieved by adding the following regularization term to the optimization objective:
@@
3.3 ToolRM Training Data Generation
To train ORMs for function-calling tasks, we require data consisting of user queries, tool catalogs, and the corresponding correct and incorrect tool calls. We construct this data by leveraging a diverse set of open-source, permissively licensed language models with function-calling capabilities. Specifically, we use publicly available function-calling datasets, which provide user queries, tool catalogs, and ground-truth tool call sequences. For each query, we prompt the models to generate tool calls using the tools specified in the dataset.
The generated tool calls are then compared against the ground-truth sequences. Outputs that deviate from the ground truth are retained as incorrect examples, while matching outputs are discarded. This procedure enables the collection of data that reflects the natural variability and error patterns of real-world models. It captures not only common mistakes but also subtle and complex failure modes that are difficult to anticipate or enumerate manually.
@@ 4 Experimental Setup
To create training data for the RM, we select open-source datasets that cover various aspects of function-calling, such as the API-Gen dataset (Liu et al., 2024c) for single-turn interactions, the Schema-Guided Dialogue (SGD) dataset (Rastogi et al., 2020) for multi-turn interactions with tool invocations and responses, and the xlam-irrelevance [^2] https://huggingface.co/datasets/MadeAgents/xlam-irrelevance-7.5k dataset for cases where the system lacks sufficient information to respond to a user query.
Since these datasets are common training datasets and our primary focus is to elicit representative incorrect behavior from the model, we follow Lin et al. (2024) and obfuscate the data samples to avoid the model regurgitating its training data. We obfuscate the samples by replacing function and parameter names with randomly generated strings and reordering the keys in the function schema.
We then use a collection of 11 permissively-licensed, open-weight models to generate the training data. The pool includes both general-purpose instruction-tuned models with function-calling capabilities and function-calling specific models, with parameter counts ranging from 0.5B to 32B. Specifically, we use the Qwen2.5-Instruct (Team, 2024) and Granite 3.3-Instruct (Granite Team, ) model series, along with Granite-20b-function-calling (Abdelaziz et al., 2024), SmolLM2 (Allal et al., 2025), Mistral-7b-Instruct-v0.3 and Mistral-Nemo-Instruct-2407.
After generating outputs from the model pool and keeping only the incorrect ones, we subsample one incorrect output per input user query to prevent over-representation from a user query in the training data. Overall, this results in 180K training data samples divided into 85K single and multi-turn data each, and 10K irrelevance data. The full list of models used to generate the training data, along with a few training data samples, is provided in Appendix A.2.
@@
We use the Qwen-2.5-Instruct (1.5B, 7B, and 14B parameter variants) models (Team, 2024; Yang et al., 2024) as the base architecture for our RMs. We initialize the RMs with the instruction-tuned model weights and replace the final language modeling head with a linear layer that maps the hidden representation to a scalar reward value.
The RMs accept the specifications of available functions, conversation history, and the generated tool call as input and produce a scalar reward as output (refer to Appendix A.3 for prompt template). We train all RMs for 1 epoch with a learning rate set to 1e-6, a cosine learning rate schedule with warmup set to 3% of total steps, and the reward centering coefficient set to 0.01.
@@
In addition to FC-RewardBench, we evaluate models on the following commonly used function-calling benchmarks: Berkeley Function Calling Leaderboard (BFCL) v3 (Patil et al., 2025), API-Bank (Li et al., 2023), ToolAlpaca (Tang et al., 2023), NexusRaven API Evaluation  [^3] https://huggingface.co/datasets/Nexusflow/NexusRaven_API_evaluation, and SealTools (Wu et al., 2024). For API-Bank, we evaluate on the Call (API-Bank-1) and Retrieval+Call (API-Bank-2) splits. Table 6 summarizes their key statistics and characteristics. We highlight that these benchmarks vary in difficulty, encompassing single and multi-turn queries, nested tool calls, and evaluation sets collected from both real users and synthetically generated.
@@
To evaluate performance on FC-RewardBench, we select eight RMs from RewardBench, spanning sizes from 3B to 70B parameters. We chose models that achieved high scores on RewardBench and support tool use in their chat template, which helps mitigate performance degradation due to prompt variability. In addition to these specialized RMs, we include six LLMs as judges, ranging from 70B to 685B parameters. See Appendix A.4 for the complete list of models.
For downstream task evaluations, we select the strongest function-calling models – the xLAM-2 series (Prabhakar et al., 2025a) – and the strongest generic instruction-tuned models – the Qwen3 series (Yang et al., 2025) – from the BFCL-v3 leaderboard. Both of these model series cover a wide range of sizes (0.6B to 70B), enabling a comprehensive assessment of ToolRM across model scales.